The product of 2 numbers is 91. The larger of the 2 numbers is one less than twice the smaller. Find the 2 numbers
Answer by TimothyLamb(4379)

x = larger number
y = smaller number
---
xy = 91
x = 2y - 1
---
xy = 91
(2y - 1)y = 91
2yy - y - 91 = 0
---
the above quadratic equation is in standard form, with a=2, b=-1 and c=-91
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
2 -1 -91
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
y = 7
y = -6.5
---
the negative root doesn't fit the problem statement, so use the positive root:
---
y = 7
x = 2y - 1
x = 2*7 - 1
x = 13
---
or
y = -6.5
x = 2y - 1
x = 2*(-6.5) - 1
x = -14
in this case x is smaller than y, and if we reverse the roles of x and y, (x=-6.5 and y=-14), the numbers don't fit the problem statement (x = 2y - 1 is not true)
---
answer:
x = 13
y = 7
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---